# Shadows of generalised Hayward spacetimes : in vacuum and with plasma

### v1

@@ V Conclusion
In this work, we have performed a detailed investigation of the shadows of various classes of spacetimes arising from different parameter ranges of the generalized Hayward metric [58]. The metric in eq.(3) depends on parameters $(\sigma,\kappa)$ and is obtained from the Hayward regular BH metric by introducing different mass parameters in the $g_{tt}$ and $g_{rr}$ components. This metric gives rise to four distinct kinds of wormholes, a regular and a singular BH solution as shown in Fig.(1) and listed in Table 1. The properties of each spacetime and their quasinormal mode frequencies have been studied in [58]. Interestingly, during the calculation of the shadow radius of these spacetimes, we observed that in the most general case, i.e. the Hayward-Damour-Solodukhin wormhole, there are two photon spheres for a specific range of $(\sigma,\kappa)$. The effective potential for this general class of wormhole exhibits single, double and triple peaks depending on the values of the metric parameters. The shadow radius for this wormhole increases as $\sigma$ increases, keeping $\kappa$ fixed while it decreases with increasing $\kappa$, with the value of $\sigma$ being fixed. The remaining wormhole and BH spacetimes are characterised by only one photon sphere.
We have studied the dependence of the shadow radius on the metric parameters for all the spacetimes and also noted their deviation from the Schwarzschild BH shadow, which occurs for $(\sigma=1,\kappa=0)$. This analysis is essential in the context of BH mimickers where certain choices of metric parameters can result in non-singular spacetimes producing shadow radii that are indistinguishable from those of a Schwarzschild BH. We have noticed that for the Hayward regular BH, the shadow radius closely matches with that of the Schwarzschild BH for small values of $\kappa$, as shown in Fig.(9(b)). Similarly, for the Damour-Solodukhin wormhole, the shadow radius approaches that of the Schwarzschild BH as $\sigma\rightarrow 1$, as seen in Fig.(4(b)). However, the shadows of both these geometries begin to deviate from the Schwarzschild shadow as $\kappa$ increases for the Hayward regular BH and as $\sigma$ becomes smaller for the Damour-Solodukhin wormhole. This is expected, as in these limits of metric parameters, the differences in the metric of these geometries become more pronounced from a Schwarzschild BH.
The above conclusions are for shadows of spacetimes when considered in vacuum i.e. in absence of any surrounding medium around the compact object. However, in real astrophysical scenarios the compact objects exist in environments. To understand the effect of such non-vacuum environment on the shadow radius, as a first step, we consider plasma surrounding the compact objects. For analytical simplicity, we focused on specific plasma profiles, namely, homogeneous $\left(\Omega(r)=k_{0}\right)$ and non-homogeneous plasma profiles $\left(\Omega(r)=\frac{k_{x}}{x}\right)$. We calculated the effects of these plasma profiles for all the classes of spacetimes and also highlighted the difference in the shadow radius in the presence and absence of plasma. The summary of the dependence of shadow radius on the metric as well as plasma parameters for all spacetimes is given in Table 2. A more realistic plasma profile cannot be handled analytically and would require a numerical approach, including other nontrivial effects such as magnetic fields and accretion, to name a few. This is well beyond this paper’s scope and will be addressed in future projects. Hence, in summary, our generalised Hayward metric provides the perfect ground for studying the shadow of various regular/singular compact objects.

### v2

@@ 6 Conclusion
In this work, we have performed a detailed investigation of the shadows cast by various spacetime geometries arising from different parameter regimes of the generalized Hayward metric [57]. The metric in Eq.(2.3) depends on parameters $(\sigma,\kappa)$ and is obtained by introducing distinct mass parameters in the $g_{tt}$ and $g_{rr}$ components of the original Hayward regular black hole metric. This metric gives rise to four distinct kinds of wormholes, along with a regular and a singular BH solution as illustrated in Fig.(1) and summarized in Table 1. While the spacetime properties and quasinormal mode spectra of these solutions have been analyzed in [57], here we focus on their shadow characteristics.
We have studied the dependence of shadow radius on the metric parameters for all spacetimes and plotted their deviation from the Schwarzschild BH shadow, corresponding to $(\sigma=1,\kappa=0)$. This analysis is especially relevant in the context of black hole mimickers, where particular parameter choices yield non-singular spacetimes whose shadow radius is nearly indistinguishable from that of a Schwarzschild BH. For instance, we observe that the Hayward regular BH yields a shadow radius nearly identical to Schwarzschild’s for small $\kappa$ values (see Fig. 9(b)). Similarly, the Damour-Solodukhin wormhole approaches the Schwarzschild shadow as $\sigma\rightarrow 1$ (Fig. 4(b)). However, significant deviations appear as $\kappa$ increases or $\sigma$ decreases, reflecting the growing difference in the underlying geometry from the Schwarzschild case. We also observe a notable feature in the most general subclass the Hayward-Damour-Solodukhin (HDS) wormhole where, for certain values of $(\sigma,\kappa)$, the effective potential admits two or three photon spheres. Depending on the metric parameters, this potential can exhibit single, double, or even triple peaks. We find that the shadow radius of the HDS wormhole increases with $\sigma$ (for fixed $\kappa$) and decreases with increasing $\kappa$ (for fixed $\sigma$). In contrast, all remaining wormhole and BH solutions possess a single photon sphere.
The above conclusions pertain to vacuum spacetimes, where no surrounding medium influences the light propagation. However, realistic astrophysical BHs exist within matter-rich environments. As a first step towards incorporating such effects, we considered the impact of plasma on the shadow structure. For analytical simplicity we focused on two plasma density profiles: a homogeneous plasma $\Omega(x)=k_{0}$ and a non-homogeneous profile $\Omega(x)=\frac{k_{x}}{x}$. We computed the shadow radii for all classes of spacetimes under these conditions and compared the results to the vacuum case. A consolidated summary of the shadow’s dependence on both the metric and plasma parameters is provided in Table 2.
Finally, we compare our results with the observational constraints from the EHT image of the shadow of Sgr $\rm A^{*}$. Upon imposing these constraints, regular black hole solutions do not see any significant change in the allowed parameter space. We also find that for wormhole solutions, the allowed parameter ranges for the spacetimes under consideration become significantly restricted. Our analysis indicates that a subclass of wormhole solutions exhibiting multi-peak effective potentials—particularly in the presence of plasma media—are more consistent with the observational data than those characterized by single-peak potentials. Interestingly, this finding stands in contrast to the conclusions drawn from quasinormal mode analyses, where single-barrier effective potentials were preferred as viable black hole mimickers [57]. If the shadow-based analysis is accurate, this tension implies that compact objects resembling wormholes might exhibit characteristic late-time echoes in their gravitational wave signals offering a complementary observational signature to be explored in future detections.Current gravitational wave detectors have limited sensitivity when it comes to detecting the ringdown from the merger of compact objects and possible echoes if present. Thus the possibility of detecting echoes through improved future gravitational wave detectors keeps the wormhole geometries characterised with multi-peak potentials a viable spacetime. Regular BH solutions also remain viable, though with a significantly narrow choice of homogeneous plasma parameter when plasma is introduced. This makes it difficult to distinguish them observationally from standard GR black holes. Future high-resolution observations, both through shadows and gravitational waves, may enhance this discriminatory power, potentially uncovering subtle features unique to exotic compact objects.
It is important to note that more realistic astrophysical scenarios involve additional complexities such as magnetic fields, accretion flows, and nontrivial plasma distributions, which cannot be treated analytically and need to be addressed. However, addressing these aspects will require a fully numerical approach and is beyond the scope of the present work. Nonetheless, the generalized Hayward metric offers a versatile framework for systematically studying the shadow characteristics of a wide spectrum of regular and singular compact objects.